# An Empirical Study and Analysis of Text-to-Image Generation Using Large Language Model-Powered Textual Representation

### v1

@@ Abstract
One critical prerequisite for faithful text-to-image generation is the accurate understanding of text inputs. Existing methods leverage the text encoder of the CLIP model to represent input prompts. However, the pre-trained CLIP model can merely encode English with a maximum token length of 77. Moreover, the model capacity of the text encoder from CLIP is relatively limited compared to Large Language Models (LLMs), which offer multilingual input, accommodate longer context, and achieve superior text representation. In this paper, we investigate LLMs as the text encoder to improve the language understanding in text-to-image generation. Unfortunately, training text-to-image generative model with LLMs from scratch demands significant computational resources and data. To this end, we introduce a three-stage training pipeline that effectively and efficiently integrates the existing text-to-image model with LLMs. Specifically, we propose a lightweight adapter that enables fast training of the text-to-image model using the textual representations from LLMs. Extensive experiments demonstrate that our model supports not only multilingual but also longer input context with superior image generation quality.
@@ 1 Introduction
To address these issues, we propose a three-stage training pipeline that efficiently integrating LLMs into existing diffusion models. The main idea of our method is a lightweight but effective adapter module to align the text features of LLMs with that of the visual-aware CLIP. In this way, LLMs could capture the visual clues contained in the input prompts, thereby drive text-to-image diffusion models to produce appropriate images. Specifically, we decompose the training procedure into three distinct stages. First, we adapt the features of LLMs into diffusion training process by aligning them with those from CLIP models, only adapter is optimized in this stage. Then, we improve the synthesis quality through end-to-end text-image training. After that, the aesthetic appeal of the generated images is enhanced by further finetuning on a carefully-curated dataset. By doing so, the textual representation capabilities of LLMs can be fully activated and the model performance is well improved in terms of text alignment, synthesis quality and image aesthetics. Notably, our model is trained with a fraction of the resources required by most text-to-image diffusion models while achieving superior synthesis quality and supporting multilingual input.
@@
  • We devise an efficient three-stage training pipeline that accomplish fast adaptation of LLM textual features with a small amount of resources, serving as an strong baseline of integrating LLMs into diffusion models and paving the way of this important topic.
  • Extensive experiments demonstrate that our model not only supports multilingual input conditions (i.e., Chinese, Japanese, Korean, etc.) but also surpasses existing text-to-image diffusion models in terms of synthesis quality, text alignment and image aesthetics.
@@ 3.2 Framework
where $h_{\text{LLM}}\in\mathbb{R}^{L\times d_{\text{LLM}}}$ and $h_{\text{adapter}}\in\mathbb{R}^{Q\times d_{q}}$, respectively denotes the output features of the LLM and the adapter. $q$ is a hyper-parameter represents the number of learnable queries, which set as $77$ in implementation.
@@
Figure 3: Three-stage training pipeline. Multilingual textual alignment enables LLM to capture the visual and textual awareness of CLIP embedding space. End-to-end text-image training fully explores the potential of LLM-derived textual features and improves generation quality. High-aesthetic finetuning further ameliorates the visual aesthetic.
@@ 3.3 Model Training Strategy
Our model is trained with three stages, namely multilingual textual alignment, end-to-end text-image training, and high-aesthetic finetuning. The overall pipeline is presented in Fig. 3. We describe details as follows.
Stage 1: Multilingual Textual Alignment. In order to introduce the text features of LLMs into the diffusion training process, we train the adapter to align the text feature of LLM with that of CLIP by merely optimizing the adapter through the proposed alignment loss. During training, we use two types of text corpora, English-only and Chinese-English paired text. For both of these datasets, the CLIP model is exclusively provided with English text, obtaining the features to be aligned, $h_{\text{clip}}$. Meanwhile, the adapter-suffixed LLM is respectively fed with English and Chinese text for English-only and Chinese-English corpora, yielding the adaption-needed LLM-derived features, $h_{\text{adapter}}$. Then, we aligning these two features by optimizing the adapter module through an alignment loss, which is defined as:
@@
where the first term constrains the magnitude of these two features should be aligned, and the second term ensures the features from the adapter-suffixed LLM align with the feature space of the CLIP model, $\lambda$ is the hyper-parameter that balance them.
Recall that the maximum token length of CLIP models is only $77$, resulting information loss on text that is longer than $77$. Therefore, directly aligning the adapter-suffixed LLM with the CLIP may bias the LLM-derived text features to short prompt text-to-image generation and hinders alignment performance. To mitigate this limitation, we adopted a segmented encoding approach. It first segments the input prompt into chunks of $77$ tokens for separate CLIP encoding, the CLIP features of each chunk are then concatenated to obtain the complete feature. Such that, the information beyond 77 tokens are well-preserved without modifying the CLIP model, allowing multilingual textual alignment on longer text data.
@@
Stage 3: High-aesthetic Fine-tuning. Stage 2 ensures the alignment of image-text embeddings of LLMs and diffusion process, but the overall image quality is relatively random and unstable. In order to guide the model to exclusively produce images with highly visual aesthetics, we further finetune the diffusion model with a carefully curated high-quality images, which have shown appealing quality gains in the community [9, 7]. Specifically, we filter $40$K exceptionally high-quality images from a large amount of filtered images, where the filtering process includes text and optical character recognition, aesthetic score filter, image-text alignment, position, and object detection, as well as human visual judgment. Further, we employ a large multimodal model [27] to caption these images for highly descriptive text conditions with specific styles. In stage 3, the model is slightly optimized for nearly $5,000$ steps with a batchsize of $64$.
Figure 4: Qualitative results of our model and competing methods. For models that do not support multilingual text conditions, we translate the given prompts into corresponding language to generate images. Our proposed method could produce images with accurate text-image alignment and higher visual quality.
@@ 4 Experiments
We conduct qualitative and quantitative experiments to validate our method. We first introduce experimental setups in Sec. 4.1. We then compare our methods with state-to-the-art models in Sec. 4.2. Finally, we show ablative studies and analyses in Sec. 4.3.
@@ 4.1 Experimental Setup
Table 1: Quantitative comparisons between our proposed method and existing baselines. Our model achieves the best or second-best results in FID and CLIP-Score, while outperforms compared methods significantly in aesthetics score (Aes).
@@
Datasets. Our three-stage training pipeline is respectively trained on pure text data (stage 1), text-image data (stage 2), and high-quality text-image data (stage 3). Regarding the text data, we use the text captions of LAION-5B [41] and collected real-world user prompts. For the paired text-image data for stage 2, we similarly employ the LAION-5B dataset and build an self-collected text-image dataset from internet, yielding a total of $43$M samples. Notably, to fully exploit LLM-powered language understanding ability as done in [4, 7], we re-caption these images with a self-trained multi-modal LLMs, providing more descriptive text conditions for training. Finally, a carefully-crafted high-quality dataset is consist of both real-world images from internet and public-available image synthesis benchmarks [46], obtaining a total of $40$K images with a average aesthetic score of $7.51$. Notably, All of the aforementioned data is filtered, pre-processed and re-captioned with a carefully designed data cleaning process.
Evaluation Metrics. Following existing philosophy of evaluating text-to-image generative models, we include three commonly-used automatic metrics, CLIP-score [14], FID [15] and aesthetics score [41] to evaluate the generated images in terms of text alignment, synthesis quality and aesthetic appeal. Considering that our model features multilingual text conditions, we gather both English and Chinese prompts for a thorough investigation. Specifically, for English-driven text-to-image generation, we use COCO [25], Flickr-30K [56], and Flickr-8K [18], and use COCO-cn [24], Flickr-8K-cn [23] for Chinese driven text-to-image generation. Note that we only utilize the text captions for producing images in inference, and Chinese-CLIP [54] is used for calculating CLIP-Score of Chinese benchmarks. The original texts are used for computing CLIP-Score, and the original images are used as reference distribution for zero-shot FID evaluation.
To better investigate the visual quality of our proposed method, we further conduct an user study to identify the human preference of different models. Presented pairs of images generated by different models, $10$ individuals are asked to select more visually appealing images. To ensure reliability of human evaluation, annotators are not aware of the image was produced by which model.
Comparison Baselines. Quantitatively, we compare our method with several existing multilingual text-to-image diffusion models, namely Alt-Diffusion [55], and Taiyi-Diffusion [49]. Moreover, for those baselines that do not support multilingual input conditions, i.e., SD-v1.5 [45] and SD-XL [32], we compare their qualitative results based on machine-translated text conditions, e.g., translating Flickr-8K-cn texts into English.
@@ 4.2 Main Results
Quantitative Results.  Tab. 1 presents the comparison results between our proposed method and existing alternatives. Owing to the aligned powerful language understanding of LLMs, our model achieves the best or the second-best FID/CLIP-Score performances under all tested benchmarks, while gains a significant improvement in visual aesthetics. It indicates that our model are able to generate images with competitive synthesis quality, textual alignment and aesthetic appeal.
@@
Qualitative Results.  Fig. 4 shows the visual outputs of our proposed method and comparison baselines, providing a clearer comparison on the synthesis quality. Clearly, the overall quality of images generated by our model is more visually appealing, and the prompts following and text-image alignment are significantly better than others. For instance, given “A brown and white dog has is playing with a white ball on the grass" as input, the image produced by our model is more realistic, the texture and visual details are more plausible. Moreover, in the last row, compared models miss generating other people in the back of the woman, and their output images look unnatural. This reflects that our model behaves better at following user prompts and synthesizing high-quality images.
Human Evaluation. In order to investigate the human preference of our model, we conduct user studies to compare with prior approaches in a pair-by-pair manner. Each time, we present users two images generated by our model and compared baselines, and then ask users to choose the better one.  Fig. 5 shows the results. It turns out that our model is consistently identified as the better model when compared to different alternatives, particularly Taiyi-Diffusion and Alt-Diffusion. Together with the above quantitative and qualitative results, the advance of our proposed method is well proved.
@@ 4.3 Ablation Study and Further Analysis
Figure 6: Multilingual outputs. Our model is capable of understanding various languages and produce corresponding image content.
@@
Multilingual Text-to-image Generation.  Fig. 6 provides the visual outputs of using various languages as the input conditions of our model. Surprisingly, our model could understand these texts well and generate images with corresponding captions. This amazing feature indicates that our model successfully integrates the powerful language understanding ability of LLMs into the text-to-image generation process, and fully exploit the potential of LLMs.
Long Prompts Generation. To testify whether our model could understand long texts, here we present the generated images under long prompts in Fig. 7. Remarkably, the proposed model could capture the meaning of prompts that are much longer than $77$ tokens and synthesize images that well align with prompts, whereas prior methods usually fail under such setting. This further reflects the powerful language understanding capability and synthesis quality of our method.
Figure 7: Long prompts driven generation results. Our model can capture long contextual text semantics and synthesize text-alignment results.
@@ 5 Conclusion
In this paper, we present an empirical study on empowering text-to-image diffusion models with the exceptional language understanding capabilities of LLMs. We suffix a lightweight adapter to LLM to enable efficient feature alignment of LLM and CLIP, enabling text-to-image diffusion models to condition on LLM-derived text features. Specifically, we design a three-stage training pipeline to fully adapt the text features of LLM to the diffusion training process. Extensive results verify the effectiveness of the proposed training pipeline, yielding LLM-powered diffusion models that are able to create images with high synthesis quality, textual alignment and aesthetic appeal. Besides, our models also support text-to-image generation under multilingual and long-prompt scenario, serving as an effective and practical baseline for bootstrapping diffusion models with LLMs.

### v2

@@ Abstract
One critical prerequisite for faithful text-to-image generation is the accurate understanding of text inputs. Existing methods leverage the text encoder of the CLIP model to represent input prompts. However, the pre-trained CLIP model can merely encode English with a maximum token length of 77. Moreover, the model capacity of the text encoder from CLIP is relatively limited compared to Large Language Models (LLMs), which offer multilingual input, accommodate longer context, and achieve superior text representation. In this paper, we investigate LLMs as the text encoder to improve the language understanding in text-to-image generation. Unfortunately, training text-to-image generative model with LLMs from scratch demands significant computational resources and data. To this end, we introduce a three-stage training pipeline, OmniDiffusion, that effectively and efficiently integrates the existing text-to-image model with LLMs. Specifically, we propose a lightweight adapter that enables fast training of the text-to-image model using the textual representations from LLMs. Extensive experiments demonstrate that our model supports not only multilingual but also longer input context with superior image generation quality. Project page: https://llm-conditioned-diffusion.github.io
@@ 1 Introduction
To address these issues, we propose a three-stage training pipeline, OmniDiffusion, that efficiently integrating LLMs into existing diffusion models. The main idea of our method is a lightweight but effective adapter module to align the text features of LLMs with that of the visual-aware CLIP. In this way, LLMs could capture the visual clues contained in the input prompts, thereby drive text-to-image diffusion models to produce appropriate images. Specifically, we decompose the training procedure into three distinct stages. First, we adapt the features of LLMs into diffusion training process by aligning them with those from CLIP models, only adapter is optimized in this stage. Then, we improve the synthesis quality through end-to-end text-image training. After that, the aesthetic appeal of the generated images is enhanced by further finetuning on a carefully-curated dataset. By doing so, the textual representation capabilities of LLMs can be fully activated and the model performance is well improved in terms of text alignment, synthesis quality and image aesthetics. Notably, our model is trained with a fraction of the resources required by most text-to-image diffusion models while achieving superior synthesis quality and supporting multilingual input.
@@
  • We devise an efficient three-stage training pipeline, OmniDiffusion, that accomplish fast adaptation of LLM textual features with a small amount of resources, serving as an strong baseline of integrating LLMs into diffusion models and paving the way of this important topic.
  • Extensive experiments demonstrate that OmniDiffusion not only supports multilingual input conditions (i.e., Chinese, Japanese, Korean, etc.) but also surpasses existing text-to-image diffusion models in terms of synthesis quality, text alignment and image aesthetics.
@@ 3.2 Framework
where $h_{\text{LLM}}\in\mathbb{R}^{L\times d_{\text{LLM}}}$ and $h_{\text{adapter}}\in\mathbb{R}^{Q\times d_{q}}$, respectively denotes the output features of the LLM and the adapter. $q$ is a hyper-parameter represents the number of learnable queries, which set as $227$ in implementation.
@@
Figure 3: Three-stage training pipeline. Multilingual textual alignment enables LLM to connect visual and textual information in the CLIP embedding space. End-to-end text-image training explores the potential of LLM-derived textual features and improves generation quality. High-aesthetic finetuning further ameliorates the visual aesthetic.
@@ 3.3 Model Training Strategy
OmniDiffusion is trained with three stages, namely multilingual textual alignment, end-to-end text-image training, and high-aesthetic finetuning. The overall pipeline is presented in Fig. 3. We describe details as follows.
Stage 1: Multilingual Textual Alignment. In order to introduce the text features of LLMs into the diffusion training process, we train the adapter to align the text feature of LLM with that of CLIP by merely optimizing the adapter through the proposed alignment loss. During training, we use two types of text corpora, English-only and Chinese-English paired text. For both of these datasets, the CLIP model is exclusively provided with English text, obtaining the features to be aligned, $h_{\text{clip}}$. Meanwhile, the adapter-suffixed LLM is respectively fed with English and Chinese text for English-only and Chinese-English corpora, yielding the adaption-needed LLM-derived features, $h_{\text{adapter}}$. Then, we align these features by optimizing the adapter module through an alignment loss, which is defined as:
@@
where the first term constrains the magnitude of these two features should be aligned, and the second term ensures the features from the adapter-suffixed LLM align with the feature space of the CLIP model, $\lambda$ is the hyper-parameter that balances them.
Recall that the maximum token length of CLIP models is only $77$, resulting information loss on text that is longer than $77$. Therefore, directly aligning the adapter-suffixed LLM with the CLIP may bias the LLM-derived text features to short prompt text-to-image generation and hinders alignment performance. To mitigate this limitation, we adopt a segmented encoding approach. It first segments the input prompt into chunks of $77$ tokens for separate CLIP encoding, the CLIP features of each chunk are then concatenated to obtain the complete feature. Such that, the information beyond 77 tokens are well-preserved without modifying the CLIP model, allowing multilingual textual alignment on longer text data.
@@
Stage 3: High-aesthetic Fine-tuning. Stage 2 ensures the alignment of image-text embeddings of LLMs and diffusion process, but the overall image quality is relatively random and unstable. In order to guide the model to exclusively produce images with highly visual aesthetics, we further finetune the diffusion model with a set of carefully curated high-quality images, which has shown appealing quality gains in the community [9, 7]. Specifically, we filter $40$K exceptionally high-quality images from a large amount of filtered images, where the filtering process includes text and optical character recognition, aesthetic score filter, image-text alignment, position, and object detection, as well as human visual judgment. Further, we employ a large multimodal model [27] to caption these images for highly descriptive text conditions with specific styles. In stage 3, the model is slightly optimized for nearly $5,000$ steps with a batchsize of $64$.
Figure 4: Qualitative results of OmniDiffusion and competing methods. For models that do not support multilingual text conditions, we translate the given prompts into corresponding language to generate images. OmniDiffusion could produce images with accurate text-image alignment and higher visual quality.
@@ 4 Experiments
We conduct qualitative and quantitative experiments to validate our method. We first introduce experimental setups in Sec. 4.1. We then compare the proposed OmniDiffusion with state-to-the-art models in Sec. 4.2. Finally, we show ablative studies and analyses in Sec. 4.3.
@@ 4.1 Experimental Setup
Table 1: Quantitative comparisons between our proposed method and existing baselines. OmniDiffusion achieves the best or second-best results in FID and CLIP-Score, while outperforms compared methods significantly in aesthetics score (Aes).
@@
Datasets. Our three-stage training pipeline is respectively trained on pure text data (stage 1), text-image data (stage 2), and high-quality text-image data (stage 3). Regarding the text data, we use the text captions of LAION-5B [41] and collected real-world user prompts. For the paired text-image data for stage 2, we similarly employ the LAION-5B dataset and build an self-collected text-image dataset from internet, yielding a total of $43$M samples. Notably, to fully exploit LLM-powered language understanding ability as done in [4, 7], we re-caption these images with a self-trained multi-modal LLMs, providing more descriptive text conditions for training. Finally, a carefully-crafted high-quality dataset is consist of both real-world images from internet and public-available image synthesis benchmarks [46], obtaining a total of $40$K images with a average aesthetic score of $7.51$.
Evaluation Metrics. Following existing philosophy of evaluating text-to-image generative models, we include three commonly-used automatic metrics, CLIP-score [14], FID [15] and aesthetics score [41] to evaluate the generated images in terms of text alignment, synthesis quality and aesthetic appeal. Considering that OmniDiffusion features multilingual text conditions, we gather both English and Chinese prompts for a thorough investigation. Specifically, for English-driven text-to-image generation, we use COCO [25], Flickr-30K [57], and Flickr-8K [18], and use COCO-cn [24], Flickr-8K-cn [23] for Chinese driven text-to-image generation. Note that we only utilize the text captions for producing images in inference, and Chinese-CLIP [55] is used for calculating CLIP-Score of Chinese benchmarks. The original texts are used for computing CLIP-Score, and the original images are used as reference distribution for zero-shot FID evaluation.
To better investigate the image visual quality of OmniDiffusion, we further conduct an user study to identify the human preference of different models. Presented pairs of images generated by different models, $10$ individuals are asked to select more visually appealing images. To ensure the reliability of human evaluation, annotators are not aware of the image was produced by which model.
Comparison Baselines. Quantitatively, we compare OmniDiffusion with several existing multilingual text-to-image diffusion models, namely Alt-Diffusion [56], and Taiyi-Diffusion [49]. Moreover, for those baselines that do not support multilingual input conditions, i.e., SD-v1.5 [45] and SD-XL [32], we compare their qualitative results based on machine-translated text conditions, e.g., translating Flickr-8K-cn texts into English.
@@ 4.2 Main Results
Quantitative Results.  Tab. 1 presents the comparison results between our proposed method and existing alternatives. Owing to the aligned powerful language understanding of LLMs, OmniDiffusion achieves the best or the second-best FID/CLIP-Score performances under all tested benchmarks while gaining a significant improvement in visual aesthetics. It indicates that our model is able to generate images with competitive synthesis quality, textual alignment and aesthetic appeal.
@@
Qualitative Results.  Fig. 4 shows the visual outputs of our proposed method and comparison baselines, providing a clearer comparison on the synthesis quality. Clearly, the overall quality of images generated by OmniDiffusion is more visually appealing, and the prompts following and text-image alignment are significantly better than others. For instance, given “A brown and white dog has is playing with a white ball on the grass" as input, the image produced by our model is more realistic, the texture and visual details are more plausible. Moreover, in the last row, compared models miss generating other people in the back of the woman, and their output images look unnatural. This reflects that OmniDiffusion behaves better at following user prompts and synthesizing high-quality images.
Human Evaluation. In order to investigate the human preference of OmniDiffusion, we conduct user studies to compare with prior approaches in a pair-by-pair manner. Each time, we present users two images generated by OmniDiffusion and compared baselines, and then ask users to choose the better one.  Fig. 5 shows the results. It turns out that our model is consistently identified as the better model when compared to different alternatives, particularly Taiyi-Diffusion and Alt-Diffusion. Together with the above quantitative and qualitative results, the advance of our proposed method is well proved.
@@ 4.3 Ablation Study and Further Analysis
Figure 6: Multilingual outputs. OmniDiffusion is capable of understanding various languages and produce corresponding image content.
@@
Multilingual Text-to-image Generation.  Fig. 6 provides the visual outputs of using various languages as the input conditions of our model. Surprisingly, OmniDiffusion could understand these texts well and generate images with corresponding captions. This amazing feature indicates that OmniDiffusion successfully integrates the powerful language understanding ability of LLMs into the text-to-image generation process, and fully exploit the potential of LLMs.
Long Prompts Generation. To testify whether our model could understand long texts, we present the generated images under long prompts in Fig. 7. Remarkably, OmniDiffusion captures the meaning of prompts that are much longer than $77$ tokens and synthesizes images that well align with prompts, whereas prior methods usually fail under such setting. This further reflects the powerful language understanding capability and synthesis quality of OmniDiffusion.
Figure 7: Long prompts driven generation results. OmniDiffusion can capture long contextual text semantics and synthesize text-alignment results.
@@ 5 Conclusion
In this paper, we present an empirical study on exploiting the Large-language-model powered textual understanding capability for text-to-image diffusion models. A lightweight but effective adapter is developed to efficiently connects LLMs’ text features with CLIP’s embeddings, making LLMs’ representation applicable for T2I tasks. Additionally, a novel three-stage training pipeline, OmniDiffusion, is designed to fully excavate LLMs’ potential and simultaneous improve the synthesis quality. Extensive results demonstrate the efficacy of each stage and the exceptional synthesis features enabled by OmniDiffusion, and our method provides an effective and practical baseline for connecting LLMs with T2I diffusion models.
@@ Appendix 0.D More Analysis Results
Analysis on Text Encoders. To quantitatively verify the advantage of LLMs over other text encoders, we trained T5-integrated and LLM-integrated diffusion models with 14,000 steps of stage 2 training, we also include T5-based Wuerstchen for a fair comparison. Considering that T5-integrated model understands English only while LLM-integrated model naively supports multilingual understanding. Furthermore, T5-integrated model performs worse than LLM-integrated in FID and CLIP-s, indicating the advantage of LLMs.
Table A3: Comparison with T5-empowered models. we trained T5-integrated and LLM-integrated diffusion models with the same training costs and procedure, we also include the popular T5-empowered diffusion model, Wuerstchen for a fair comparison.
Models Flickr8k-en COCO-en
FID CLIP-s Aes FID CLIP-s Aes
Ours (s2 T5) 73.75 25.11 5.93 16.20 24.16 5.78
Ours (s2 LLM) 72.72 26.23 5.97 15.50 24.92 5.87
Wuerstchen (T5) 75.81 27.24 5.82 22.96 25.90 5.75
Analysis on Training and Inference Costs. As for training costs, our model benefits from the proposed three-stages training paradigm that it only incorporates a lightweight adapter and is trained to efficiently connect LLMs with diffusion models. Compared with the training complexity and resource requirements of training LLM-empowered diffusion models from scratch, our model achieves a better trade-off and is highly democratizing, serving as a strong baseline to deliver multilingual and long-prompts text understanding for T2I tasks. As for the inference costs, we conduct experiments on diffusion models integrated with different text encoders and summarize the results in Tab. A4. We can conclude that the main inference costs come from the inverse diffusion process, as the computational costs of integrating LLMs have minimal effect on the inference time.
Table A4: Inference cost comparison with other text encoders. we include CLIP, T5, and a popular bilingual LLM, Baichuan2-7B, for comparison. In this table, we demonstrate the inference time under different text encoders and inference steps. As can be concluded, the inference time is highly consistent with the number of inference steps while showing little consistency with the employed text encoder.
Text Encoders Inference Steps
25 50 100
CLIP 3.13 6.04 11.84
T5 3.20 6.13 12.09
Baichuan2-7B 3.21 6.16 12.03
Analysis on Stage 1 Training. The average token length of the data used in stage 1 is 40.2, 11.7% of which exceed 77. During stage 1 training, we preserve the information beyond 77 tokens by applying a segmented encoding approach, thereby allowing multilingual textual alignment on longer text data. In stage 1 training, the model also supports other languages, such as Japanese. It is well-known in the community that, with well-pretrained multilingual embeddings, models align with one specific language also partially align with other languages. Consequently, the model is trained on English and Chinese corpora in stage 1 training but preserved its multilingual understanding abilities. However, to enhance model performance on other languages, a further stage 1 training on these language would be significantly beneficial.
Analysis on Comparison with Other LLM-empowered Diffusion Models. There are some works that also explore how to enhance diffusion models using LLMs, such as ParaDiffusion [52] and BiDiffusers [61], which bear resemblance to us. Here we clarify that this paper bootstraps text-to-image diffusion models with LLM-derived text features via suffixing a transformer-based adapter into LLMs. Moreover, a novel and efficient three-stage training scheme to align LLMs with a pretrained diffusion model, enabling high-quality and high-aesthetic, as well as appealing prompt-following image synthesis, especially multilingual and long-prompts generation. Notably, our work provides comprehensive experiments and analysis, namely automatic evaluation, user study, case study, and ablation study, serving as an valuable empirical study and strong baseline for LLM-integrated T2I diffusion models.
Further, our method differs Paradiffusion and BiDiffuser in the following aspects. On the one hand, ParaDiffusion does not involve a textual alignment stage for faster convergence and is trained on almost 6.98 $\times$ of the data used to train our model. Besides, ParaDiffusion use projection layer to connect a LoRA finetuned LLM with the diffusion models while we use a transformer adapter modules to connect a frozen LLM with the diffusion models. On the other hand, BiDiffuser is constrained by its modified diffusion models and its training pipeline cannot be applied to other off-the-shelf diffusion pipelines such as SDXL and Pixart, while our methods do not require any modification to diffusion pipeline, which can be applied to any diffusion pipelines, even BiDiffuser. Importantly, our empirical findings and practical solution (See Fig. A1) might inspire more interesting works to connect LLMs with text-to-image diffusion models in the community.
Figure A1: Comparison with other LLM empowered diffusion models. BiDiffuser is not included because of the official released code is not executable. As can be seen, compared with ParaDiffusion, our model synthesizes images with higher quality, fidelity, aesthetic.